# Isomorphism classes of generating sets

### v1

@@ 5. Applications to cofinal types and cardinal characteristics
$(1),(2a),(2b)$ follows from Proposition 5.10 and Corollaries 3.12,4.7, and 4.9 respectively. ∎
Gitik notified us that he was able to obtain $(2b)$ independently for the depth spectrum of the club filter at a measurable cardinal from optimal assumptions. Next, let us consider the Tukey-related notion of cohesiveness due to Kanamori [31]. This notion is also known as Galvin’s property [6, 9, 7, 2]:
@@
As mentioned in the paragraph following Definition 5.12, not being $(\lambda,\lambda)$-cohesive is equivalent to $\lambda\in\operatorname{Sp}_{T}(U)$. Hence the first part follows from Proposition 5.10. For the second part, let us prove that $U$ is $(\lambda_{0},\lambda_{1})$-cohesive. Fix ${\langle}X_{i}\mid i<\lambda_{1}{\rangle}\subseteq U$. We need to find $\lambda_{0}$-many sets whose intersection is in $U$. Let $\mathcal{B}=(b_{i,j})_{(i,j)\in\lambda_{0}\times\lambda_{1}}$ be a base for $U$ witnessing that $U$ is a simple $P_{\lambda_{0}\times\lambda_{1}}$-point. For each $i<\lambda_{1}$ there is $\beta_{i}<\lambda_{1}$ and $\alpha_{i}<\lambda_{0}$ such that $b_{\alpha_{i},\beta_{i}}\subseteq^{*}X_{i}$. There is $I\in[\lambda_{1}]^{\lambda_{1}}$ and $\alpha^{*}<\lambda_{0}$ such that for every $i\in I$, $\alpha_{i}=\alpha^{*}$. Consider the first $\lambda_{0}$-many indices $\{\beta_{i_{\gamma}}\mid\gamma<\lambda_{0}\}\subseteq I$. Let $\beta^{*}=\sup_{\gamma<\lambda_{0}}\beta_{i_{\gamma}}+1<\lambda_{1}$. Then $b_{\alpha^{*},\beta^{*}}\subseteq^{*}b_{\alpha^{*},i_{\gamma}}\subseteq^{*}X_{i_{\gamma}}$ for every $\gamma<\lambda_{0}$. Since $\lambda_{0}>\kappa$ is regular, we can find $J\subseteq\{i_{\gamma}\mid\gamma<\lambda_{0}\}$ still of size $\lambda_{0}$ and some $\xi<\kappa$ such that $b_{\alpha^{*},\beta^{*}}\setminus\xi\subseteq\bigcap_{j\in J}X_{j}$, as wanted. ∎
@@
Relative to a supercompact cardinal, there is a supercompact cardinal $\kappa$ such that for any regular cardinals
$\kappa<\lambda_{0}<\lambda_{1}<...<\lambda_{n}<\lambda_{n+1},$
there is a $\kappa$-directed closed, $\kappa^{+}$-cc generic extension with a simple $P_{\lambda_{0}\times...\times\lambda_{n}}$-point ultrafilter and $2^{\kappa}=\lambda_{n+1}$.
@@
Let $S$, $\{X_{a}\mid a\in\lambda_{0}\times...\times\lambda_{n-1}\times S\}$, and $U$ be as in the proof of Theorem 4.2, where $S\subseteq\lambda_{n+1}^{+}$ is a stationary set, and note that we skipped $\lambda_{n}$. By the items before Claim 4.3, and by shrinking $S$ if necessary, we may assume that the limit $s^{*}$ of the first $\lambda_{n+1}+\lambda_{n}$-many points of $S$ satisfies that:
$U^{*}=U\cap V[G_{\prod_{i<n}\lambda_{i}\times s^{*}}]\in V[G_{\prod_{i<n}\lambda_{i}\times s^{*}}]=V^{*}$
We claim that $U^{*}$ and $V^{*}$ are as wanted. Clearly, in $V^{*}$, $2^{\kappa}=\lambda_{n+1}$. Also $U^{*}$ is an $V^{*}$-ultrafilter on $\kappa$, which is generated by $\{X_{a}\mid a\in\lambda_{0}\times...\times\lambda_{n-1}\times(S\cap s^{*})\}$ hence $U^{*}$ is a simple $P_{\lambda_{0}\times...\times\lambda_{n-1}\times(S\cap s^{*})}$-point. Since $\lambda_{0}\times...\times\lambda_{n-1}\times\lambda_{n}$ cofinally embeds into $\lambda_{0}\times...\times\lambda_{n-1}\times S\cap s^{*}$, by Proposition 2.1, $U^{*}$ is a simple $P_{\lambda_{0}\times...\times\lambda_{n-1}\times\lambda_{n}}$-point. ∎
@@
The existence of a simple $P_{\mathbb{D}}$-point ultrafilter $U$ poses restrictions on cardinal characteristics. For example, by [10] on a measurable cardinal $\kappa$, $\min(\operatorname{Sp}_{T}(\mathbb{D}))=\min(\operatorname{Sp}_{T}(U))\leq\mathfrak{b}_{\kappa}$ and $\mathfrak{ch}(\mathbb{D})=\mathfrak{ch}(U)\geq\mathfrak{d}_{\kappa}$. However, it is unclear if this equality must hold or if there are any other limitations (See Questions 6.7-6.9).
@@ 6. Problems
It is consistent to have a simple $P_{\lambda_{0}\times\lambda_{1}\times\lambda_{2}}$-point ultrafilter on $\omega$ for some regular cardinals $\omega<\lambda_{0}<\lambda_{1}<\lambda_{2}$?
More generally we ask:
@@
What is the class of countably directed well-founded partial orders $\mathbb{D}$ that can be realized as a simple $P_{\mathbb{D}}$-point ultrafilter over $\omega$? over a measurable cardinal?
Question 6.3.
Question 6.4.
@@
In this paper, we focused on $\subseteq^{*}$. Still, one might also be interested in the order $\supseteq$ on an ultrafilter $U$, especially since it is connected to the Tukey-type of $(U,\supseteq)$.
Question 6.5.
@@
On $\omega$, [14] showed that it is consistent to have a simple $P_{\aleph_{1}}$-point and a simple $P_{\aleph_{2}}$-point. Nykos [37] proved that if there is a simple $P_{\lambda}$-point on $\omega$ then either $\lambda=\mathfrak{b}_{\kappa}$ or $\lambda=\mathfrak{d}_{\kappa}$. The following questions relate to generalizations of these facts:
Question 6.9.

### v2

@@ 5. Applications to cofinal types and cardinal characteristics
Using Proposition 5.13, $(1),(2a),(2b)$ follow from Theorem 4.11, 4.6, and Corollary 4.10 respectively. ∎
Gitik [27] notified us that he was able to obtain $(2b)$ independently for the depth spectrum of the club filter at a measurable cardinal from optimal assumptions. Next, let us consider the Tukey-related notion of cohesiveness due to Kanamori [31]. This notion is also known as Galvin’s property [3, 6, 4, 11]:
@@
As mentioned in the paragraph following Definition 5.15, not being $(\lambda,\lambda)$-cohesive is equivalent to $\lambda\in\operatorname{Sp}_{T}(U)$. Hence the first part follows from Proposition 5.13. For the second part, let us prove that $U$ is $(\lambda_{0},\lambda_{1})$-cohesive. Fix ${\langle}X_{i}\mid i<\lambda_{1}{\rangle}\subseteq U$. We need to find $\lambda_{0}$-many sets whose intersection is in $U$. Let $\mathcal{B}=(B_{i,j})_{(i,j)\in\lambda_{0}\times\lambda_{1}}$ be a base for $U$ witnessing that $U$ is a simple $P_{\lambda_{0}\times\lambda_{1}}$-point. For each $i<\lambda_{1}$ there is $\beta_{i}<\lambda_{1}$ and $\alpha_{i}<\lambda_{0}$ such that $B_{\alpha_{i},\beta_{i}}\subseteq^{*}X_{i}$. There is $I\in[\lambda_{1}]^{\lambda_{1}}$ and $\alpha^{*}<\lambda_{0}$ such that for every $i\in I$, $\alpha_{i}=\alpha^{*}$. Consider the first $\lambda_{0}$-many indices $\{i_{\gamma}\mid\gamma<\lambda_{0}\}\subseteq I$. Let $i^{*}=\sup_{\gamma<\lambda_{0}}i_{\gamma}+1<\lambda_{1}$. Then $B_{\alpha^{*},i^{*}}\subseteq^{*}B_{\alpha^{*},i_{\gamma}}\subseteq^{*}X_{i_{\gamma}}$ for every $\gamma<\lambda_{0}$. Since $\lambda_{0}>\kappa$ is regular, we can find $J\subseteq\{i_{\gamma}\mid\gamma<\lambda_{0}\}$ still of size $\lambda_{0}$ and some $\xi<\kappa$ such that $B_{\alpha^{*},i^{*}}\setminus\xi\subseteq\bigcap_{j\in J}X_{j}$, as wanted. ∎
@@
Suppose that $\kappa$ is a supercompact cardinal and let $\mathbb{D}$ be a $\kappa^{+}$-directed well-founded set with $|\mathbb{D}|\leq\lambda$. Then there is a ($\kappa$-$c$ $\ast\ \lambda^{+}$-closed $\ast\ \kappa^{+}$-cc) generic extension with a simple $P_{\mathbb{D}}$-point ultrafilter and $2^{\kappa}=\lambda$.
@@
Let $S^{*}_{\dot{U}}\subseteq\lambda^{+}$, $(B_{a})_{a\in S^{*}_{\dot{U}}\times\mathbb{D}}$, and $U$ be as in Theorem 4.7. Arguing as in the proof 4.6, we choose $s^{*}\in S^{*}_{\dot{U}}$ a limit point of $S^{*}_{\dot{U}}$ with an appropriate cofinal subset included in $S^{*}_{\dot{U}}$. This time we make sure that $s^{*}\geq\lambda$. Let
$U^{*}=U\cap V[G_{s^{*}\times\mathbb{D}}]\in V[G_{s^{*}\times\mathbb{D}}]=V^{*}$
The ultrafilter $U^{*}$ is a simple $P_{\mathbb{D}}$-point. Also, since $s^{*}\geq\lambda$, in $V^{*}$, $2^{\kappa}=\lambda$, as wanted. ∎
@@
The existence of a simple $P_{\mathbb{D}}$-point ultrafilter $U$ poses restrictions on cardinal characteristics. For example, by [7] on a measurable cardinal $\kappa$, $\min(\operatorname{Sp}_{T}(\mathbb{D}))=\min(\operatorname{Sp}_{T}(U))\leq\mathfrak{b}_{\kappa}$ and $\mathfrak{ch}(\mathbb{D})=\mathfrak{ch}(U)\geq\mathfrak{d}_{\kappa}$. However, it is unclear if this equality must hold or if there are any other limitations (See Questions 6.5-6.7).
@@ 6. Problems
In this paper, we focused on $\subseteq^{*}$. Still, one might also be interested in the order $\supseteq$ on an ultrafilter $U$, especially when considering the Tukey-type of $(U,\supseteq)$.
Question 6.3.
@@
On $\omega$, [14] showed that it is consistent to have a simple $P_{\aleph_{1}}$-point and a simple $P_{\aleph_{2}}$-point. Nyikos [39] proved that if there is a simple $P_{\lambda}$-point on $\omega$ then either $\lambda=\mathfrak{b}$ or $\lambda=\mathfrak{d}$. The following questions relate to generalizations of these facts:
Question 6.7.